# Does basis set superposition error significantly affect post-CCSD(T) corrections?

### v2

@@ I Introduction
Now it is indeed true that full counterpoise does not guarantee hewing closer to the CBS limit: as shown by Burns et al.[2] for orbital WFT calculations, and by Brauer et al.[3] for explicitly correlated F12 calculations, error compensation may take place between BSSE (which always overbinds) and IBSI (intrinsic basis set incompleteness, which almost invariably underbinds). Hence for small basis sets, complete neglect of BSSE may actually be beneficial, and for medium-size basis sets, ‘half-counterpoise’ (average of corrected and uncorrected interaction energies) tends to offer superior performance.[4, 2, 3]
@@
In computational thermochemistry, however, the IBSI overwhelms BSSE to such an extent that most researchers make no effort to apply BSSE corrections. This is particularly the case for total atomization energies (TAE), which are the quantum chemical ‘cognates’ of heats of formation $\Delta H^{\circ}_{f}$. Besides, higher-accuracy computational thermochemistry protocols like ccCA[5, 6] by the Wilson group, HEAT by the Stanton group,[7, 8, 9, 10] Weizmann-$n$ by our own group,[11, 12, 13] and the FPD (Feller-Peterson-Dixon, see Refs.14, 15 and references therein) all entail some variant of basis set extrapolation. If the latter works properly, it ought to eliminate BSSE altogether. (Indeed, we recently[16] exploited this fact to ‘reverse-engineer’ basis set extrapolations.)
@@
Studies of noncovalent interactions, with rare exceptions (such as Refs.17, 18, 19) stick to the CCSD(T)[20, 21] ‘gold standard of quantum chemistry’ and ignore post-CCSD(T) corrections. However, in thermochemistry, especially for TAEs, it is well-known (see Refs.[22, 7] for early reports) that kJ/mol accuracy cannot be achieved without them. CCSD(T) in fact outperforms the more rigorous CCSDT owing to a well-established error compensation (e.g., Refs.7, 9, 22, 11): higher-order triples, $T_{3}-(T)$, are almost always antibonding, while connected quadruples (Q) are universally bonding.
@@
On the one hand, CCSDT(Q) and especially CCSDTQ on the one hand have very steep CPU time scalings of $O(n_{\rm occ.}^{4}N_{\rm virt.}^{5})$ and $O(n_{\rm occ.}^{4}N_{\rm virt.}^{6})$, respectively. On the other hand, these higher-order corrections converge much more rapidly with the basis set than the overall correlation energy.[12] Consequently, post-CCSD(T) corrections tend to be evaluated in very small basis sets, and subsequently applied additively. For instance, in W4 theory,[11] the $T_{3}-(T)$ correction is extrapolated from cc-pVDZ and cc-pVTZ basis sets, the (Q) term in a cc-pVTZ basis set, and $T_{4}-(Q)$ in just a cc-pVDZ basis set. (In W4lite theory, just CCSDT(Q)/cc-pVDZ is done for the quadruples.)
@@
This then leads us to the main research question of the present paper: are such corrections materially affected by BSSE corrections? The issue was raised by a reviewer of Ref.19, where we showed that post-CCSD(T) contributions for cohesive energies of water clusters approach 1 kcal/mol for isomers of (H2O)20.
@@ II Computational Details
The CCSDT(Q),[23] CCSDT,[24] CCSD(T)Λ,[25, 26, 27, 28] and CCSD(T) calculations reported in this work were carried out using a combination of the MOLPRO 2024.1,[29] CFOUR,[30], and MRCC[31] electronic structure program systems, run on the CHEMFARM cluster of the Faculty of Chemistry at Weizmann. Owing to issues with inconsistent UHF solutions in the presence of ghost atoms, many of the small-molecule counterpoise data were generated using the MRCC interfaces of MOLPRO or CFOUR. For the noncovalent interactions, only closed-shell species are involved, and hence these calculations were carried out using standalone CFOUR (as memory permitted) or MRCC.
@@
For dissociation energies, $D_{e}$, of diatomic molecules, we applied the standard Boys-Bernardi counterpoise definition. For the TAEe values (total atomization energies) of small polyatomics, we used the SSFC (site-site function counterpoise) generalization of Wells and Wilson[36], which effectively amounts to evaluating each atomic energy in the full molecular basis set.
@@ III.1.1 Initial check for diatomic molecules
(a) Extrapolation exponents from Table 5 of Ref.41
@@
(a) Extrapolation exponents from Table 5 of Ref.41. ClOO excluded.
@@
The situation for higher-order connected triples $T_{3}-(T)$, however, is somewhat different. For the cc-pVDZ basis set, we find 0.043 kcal/mol (i.e., 0.18 kJ/mol), which however drops to 0.026 kcal/mol when diffuse functions are added, and to 0.013 kcal/mol when we move things one notch up to cc-pVTZ. Both W4 and HEAT apply cc-pV{D,T}Z extrapolations to the higher-order triples. If we do so here (with extrapolation parameters taken from table V in Ref.41), the RMS BSSE is just 0.008 kcal/mol, which may be justifiable to neglect in view of other, larger sources of uncertainty such as residual basis set incompleteness in the CCSD(T) component.[13] In fact, for the cc-pV{T,Q}Z basis sets used in W4.3 theory, the BSSE will be even more negligible.
@@ III.2.1 Small noncovalent dimers: the A24x8 dataset
Noncovalent interactions are very different in their behavior from atomization energies, in that for most noncovalent complexes, in that MP2 is already a decent to good starting point (except for $\pi$-stacking and related). Thus, CCSD-MP2 and (T) are commonly evaluated using relatively small basis sets (see, e.g., Refs.34, 42 and references therein).
@@ III.2.2 Not-so-small noncovalent complexes: the S66 dataset
CCSDT – CCSD(T) 0.008 64a 0.013 64
CCSDT – CCSD(T)Λ 0.007 64 0.006 64
CCSDT – CCSDT-3 0.007 64 0.007 64
CCSDT(Q) – CCSDT 0.007 59 0.008 24
CCSDT(Q) – CCSDT-3 0.011 59 0.014 24
CCSDT(Q) – CCSD(T) 0.003 59 0.005 24
@@
For this dataset, we will alas have to limit ourselves to the cc-pVDZ(no p on H), a.k.a., cc-pVDZ(d,s), basis set. We were able to obtain full CCSDT BSSE corrections for 61 out of 66 systems, and CCSDT(Q) for about two dozen. (It bears reiterating that, while the dimers often posed memory or computation time requirements that exceeded our available resources, the evaluation of counterpoise corrections does not require the dimers AB, only the monomers in the full dimer basis set A(B) and B(A), as well as naturally the monomers in their own basis set.)
@@ IV Conclusions
Acknowledgement
@@
Data availability statement
Raw data can be obtained from the corresponding author upon reasonable request.

### v3

@@ I Introduction
Now it is indeed true that full counterpoise does not guarantee hewing closer to the CBS limit: as shown by Burns et al.[2] for orbital WFT calculations, and by Brauer et al.[3] for explicitly correlated[4, 5, 6] F12 calculations, error compensation may take place between BSSE (which always overbinds) and IBSI (intrinsic basis set incompleteness, which almost invariably underbinds). Hence for small basis sets, complete neglect of BSSE may actually be beneficial, and for medium-size basis sets, ‘half-counterpoise’ (average of corrected and uncorrected interaction energies) tends to offer superior performance.[7, 2, 3]
@@
In computational thermochemistry, however, the IBSI overwhelms BSSE to such an extent that most researchers make no effort to apply BSSE corrections. This is particularly the case for total atomization energies (TAE), which are the quantum chemical ‘cognates’ of heats of formation $\Delta H^{\circ}_{f}$. The ‘raw’ and CP-corrected TAEs are defined analogously to the NCI situation as (for example, for a triatomic):
${\rm TAE}_{\rm raw}[{\rm ABC}]$ $=$ $E[{\rm A}]+E[{\rm B}]+E[{\rm C}]-E[{\rm ABC}]$
${\rm TAE}_{\rm CP}[{\rm ABC}]$ $=$ $E[{\rm A(BC)}]+E[{\rm(A)B(C)}]+E[{\rm(AB)C}]-E[{\rm ABC}]$
$\Delta_{\rm CP}{\rm TAE}_{\rm CP}[{\rm ABC}]$ $=$ $E[{\rm A}]+E[{\rm B}]+E[{\rm C}]$ (2)
$-E[{\rm A(BC)}]-E[{\rm(A)B(C)}]-E[{\rm(AB)C}]$
$=$ ${\rm TAE}_{\rm raw}[{\rm ABC}]-{\rm TAE}_{\rm CP}[{\rm ABC}]$
where the Wells-Wilson[8] SSFC (site-site function counterpoise) $n$-body generalization of Eq.(1) has been applied. (This effectively amounts to evaluating each atomic energy in the full molecular basis set.)
Higher-accuracy computational thermochemistry protocols like ccCA[9, 10] by the Wilson group, HEAT by the Stanton group,[11, 12, 13, 14] Weizmann-$n$ by our own group,[15, 16, 17] and the FPD (Feller-Peterson-Dixon, see Refs.18, 19 and references therein) all entail some variant of basis set extrapolation. If the latter works properly, it ought to eliminate BSSE altogether. (Indeed, we recently[20] exploited this fact to ‘reverse-engineer’ basis set extrapolations.)
@@
Studies of noncovalent interactions, with rare exceptions (such as Refs.21, 22, 23) stick to the CCSD(T)[24, 25] ‘gold standard of quantum chemistry’ and ignore post-CCSD(T) corrections. However, in thermochemistry, especially for TAEs, it is well-known (see Refs.[26, 11] for early reports) that kJ/mol accuracy cannot be achieved without them. CCSD(T) in fact outperforms the more rigorous CCSDT owing to a well-established error compensation (e.g., Refs.11, 13, 26, 15): higher-order triples, $T_{3}-(T)$, are almost always antibonding, while connected quadruples (Q) are universally bonding.
On the one hand, CCSDT(Q) and especially CCSDTQ have very steep CPU time scalings of $O(n_{\rm occ.}^{4}N_{\rm virt.}^{5})$ and $O(n_{\rm occ.}^{4}N_{\rm virt.}^{6})$, respectively. On the other hand, these higher-order corrections converge much more rapidly with the basis set than the overall correlation energy.[16] In response to a reviewer comment, we offer Table 1 as an illustration, compiled from data in the Supporting Information of Ref.[27]. The statistics given there cover a 65-molecule subset of the 200-molecule W4-17 thermochemical benchmark;[28] the subset spans a broad range of nondynamical (static) correlation character, from essentially pure dynamical correlation in H2O and CH4 at one end, to strong static correlation in O3, singlet C2, and BN at the other end. It is clearly seen in Table 1 that RMS contributions taper off rapidly as the connected excitation level increases, to reach insignificance beyond CCSDTQ(5)Λ. In tandem, it is also seen that basis set convergence becomes ever more rapid, with even unpolarized cc-pVDZ(p,s) yielding surprisingly small errors beyond CCSDT(Q), and ultimately dwindling down into numerical noise.
Table 1: Basis set convergence of RMS post-CCSD(T) contributions to the TAEs (kcal/mol) of the 65-molecule W4.3 subset of the W4-17 thermochemical benchmark.[28]
$T_{3}-(T)$ $(Q)$ $T_{4}-(Q)$ $(Q)_{\Lambda}-(Q)$ $T_{4}-(Q)_{\Lambda}$ $(5)_{\Lambda}$ $T_{5}-(5)_{\Lambda}$ $(6)_{\Lambda}$ $T_{6}-(6)_{\Lambda}$ $T_{7}$
nihil 0.84 1.29 0.26 0.17 0.11 0.08 0.015 0.009 0.001 0.001
cc-pVDZ(p,s) 0.53 0.41 0.09 0.07 0.06 0.02 0.003 0.002 0.001 REF
cc-pVDZ(d,s) 0.54 0.25 0.03 0.03 0.01 0.01 REF REF REF
cc-pVTZ(f,p) 0.18 0.11 0.012 0.006 0.007 REF
cc-pVQZ(g,d) 0.09 0.04 REF REF REF
cc-pV5Z(h,f) 0.05 0.02
cc-pV{Q,5}Z REF REF
The underlying data were extracted from the ESI of Ref.27, except for the connected sextuples with the cc-pVDZ(d,s) basis set and septuples with the cc-pVDZ(p,s) basis set, which were calculated for the present paper using the general coupled cluster implementation[29, 30] in MRCC;[31] raw energies can be found in the present paper’s ESI.
@@
Consequently, post-CCSD(T) corrections tend to be evaluated in very small basis sets, and subsequently applied additively. For instance, in W4 theory,[15] the $T_{3}-(T)$ correction is extrapolated from cc-pVDZ and cc-pVTZ basis sets (commonly indicated by the shorthand cc-pV{D,T}Z)), the (Q) term in a cc-pVTZ basis set, and $T_{4}-(Q)$ in just a cc-pVDZ basis set. (In W4lite theory, just CCSDT(Q)/cc-pVDZ is done for the quadruples.)
@@
This then leads us to the main research question of the present paper: are such corrections materially affected by BSSE corrections? The issue was raised by a reviewer of Ref.23, where we showed that post-CCSD(T) contributions for cohesive energies of water clusters approach 1 kcal/mol for isomers of (H2O)20.
@@ II Computational Details
The CCSDT(Q),[32] CCSDT,[33] CCSD(T)Λ,[34, 35, 36, 37] and CCSD(T) calculations reported in this work were carried out using a combination of the MOLPRO 2024.1,[38] CFOUR,[39], and MRCC[31] electronic structure program systems, run on the CHEMFARM cluster of the Faculty of Chemistry at Weizmann. Owing to issues with inconsistent UHF solutions in the presence of ghost atoms, many of the small-molecule counterpoise data were generated using the MRCC interfaces of MOLPRO or CFOUR. For the noncovalent interactions, only closed-shell species are involved, and hence these calculations were carried out using standalone CFOUR (as memory permitted) or MRCC. The latter code was likewise used for some of the additional data in Table 1, using the algorithms presented in Refs.[29, 30].
@@
For dissociation energies, $D_{e}$, of diatomic molecules, we applied the standard Boys-Bernardi counterpoise definition. For the TAEe values (total atomization energies) of small polyatomics, we used the SSFC (site-site function counterpoise) generalization of Wells and Wilson[8].
@@ III.1.1 Initial check for diatomic molecules
(a) Extrapolation exponents from Table 5 of Ref.47
@@
(a) Extrapolation exponents from Table 5 of Ref.47. ClOO excluded.
@@
The situation for higher-order connected triples $T_{3}-(T)$, however, is somewhat different. For the cc-pVDZ basis set, we find 0.043 kcal/mol (i.e., 0.18 kJ/mol), which however drops to 0.026 kcal/mol when diffuse functions are added, and to 0.013 kcal/mol when we move things one notch up to cc-pVTZ. Both W4 and HEAT apply cc-pV{D,T}Z extrapolations to the higher-order triples. If we do so here (with extrapolation parameters taken from table V in Ref.47), the RMS BSSE is just 0.008 kcal/mol, which may be justifiable to neglect in view of other, larger sources of uncertainty such as residual basis set incompleteness in the CCSD(T) component.[17] In fact, for the cc-pV{T,Q}Z basis sets used in W4.3 theory, the BSSE will be even more negligible.
@@ III.2.1 Small noncovalent dimers: the A24x8 dataset
Noncovalent interactions are very different in their behavior from atomization energies, in that for most noncovalent complexes, in that MP2 is already a decent to good starting point (except for $\pi$-stacking and related). Thus, CCSD-MP2 and (T) are commonly evaluated using relatively small basis sets (see, e.g., Refs.41, 48 and references therein).
@@ III.2.2 Not-so-small noncovalent complexes: the S66 dataset
CCSDT – CCSD(T) 0.008 66 0.013 64a
CCSDT – CCSD(T)Λ 0.007 66 0.006 64
CCSDT – CCSDT-3 0.007 66 0.007 64
CCSDT(Q) – CCSDT 0.007 59 0.008 25
CCSDT(Q) – CCSDT-3 0.011 59 0.014 25
CCSDT(Q) – CCSD(T) 0.003 59 0.005 25
@@
For this dataset, we will alas have to limit ourselves to the cc-pVDZ(no p on H), a.k.a., cc-pVDZ(d,s), basis set. We were able to obtain full CCSDT BSSE corrections for 64 out of 66 systems, and CCSDT(Q) for about two dozen. (It bears reiterating that, while the dimers often posed memory or computation time requirements that exceeded our available resources, the evaluation of counterpoise corrections does not require the dimers AB, only the monomers in the full dimer basis set A(B) and B(A), as well as naturally the monomers in their own basis set.)
@@ IV Conclusions
Author contributions
MS brought on the subject matter. VF performed the calculations and curated the data for the TAE BSSE corrections. ES carried out the A24x8 calculations and curated the relevant data. ES and JMLM shared the S66 calculations. JMLM oversaw the project and wrote the first draft, while all authors contributed to the completed manuscript.
Acknowledgments
@@
Financial disclosure
None reported.
Conflict of interest
The authors declare no potential conflicts of interest.
Supporting information
Five spreadsheets in Microsoft Excel format containing the raw data and workup for each table; Cartesian coordinates of the A24x8 species. Additional raw data may be obtained from the corresponding author upon reasonable request.
Table of Contents Graphic
(at magnification 200%)
Accompanying text for Table of Contents Graphic: